# Time-dependent Regularized 13-Moment Equations with Onsager Boundary Conditions in the Linear Regime

### v2

@@ 1. Introduction
Instead of directly solving the Boltzmann equation, many researchers focus on the macroscopic moments of the distribution function and modify classical fluid equations, such as the Navier-Stokes equations, to describe the mechanics of moderately rarefied gases. Two prominent techniques used to derive higher-order (in terms of Knudsen number) macroscopic fluid equations from the Boltzmann equation are the Grad’s moment method in [17] and the Chapman-Enskog expansion in [12]. Despite their successes in capturing near-equilibrium effects, the Grad’s 13-moment equations may experience a loss of hyperbolicity and unphysical subshocks. Moreover, the second-order Burnett and third-order super-Burnett equations, which are derived from the Chapman-Enskog expansion, suffer from instability. Recent researches address these limitations by employing various regularization techniques ([21, 24, 34, 6, 26, 18]) to enhance the accuracy and stability of the methods across a broader range of rarefied gas dynamics.
@@
In this work, we will investigate the regularized version of Grad’s 13-moment equations proposed in [34], which can be interpreted as a Chapman–Enskog like procedure applied to the Grad’s 13 moments. This regularization, abbreviated as R13, has been developed for both Maxwell ([32]) and non-Maxwell molecules ([36, 11]) and validated across diverse applications ([37, 28, 39]). The R13 equations are attractive since the equations include only second-order derivatives while attaining the super-Burnett order. Here the super-Burnett order means that the super-Burnett equations, which are the result of the Chapman-Enskog expansion up to the third order, can be derived from the R13 equations. Note that the original super-Burnett equations contain fourth-order derivatives, causing significant difficulties in the numerical discretization and formulation of boundary conditions. Due to the contribution of higher-order terms, the R13 equations can describe various rarefaction effects such as Knudsen boundary layers and heat fluxes from cold areas to hot areas ([27]). Also, the R13 equations are equipped with reliable boundary conditions ([41, 33]) and have been shown to be robust in numerical simulations ([27]). Recently in [40], Onsager boundary conditions are proposed to ensure the stability of boundary value problems of linear R13 equations for Maxwell molecules. However, it is not straightforward to generalize these Onsager boundary conditions to non-Maxwell molecules, even in the linear case. The main reason is the loss of a symmetric structure required by the Onsager boundary conditions (see [7]). To remedy this, in [10], the linear R13 equations for arbitrary elastic collision models were re-derived using a different method. The new derivation relies on the simpler form of the Chapman-Enskog expansion for steady-state equations compared with the time-dependent case, so that it cannot be applied directly to dynamic problems.
@@
The rest of this paper is organized as follows. Section 2 introduces the variables and then shows our main results including the explicit expressions of linear R13 equations with boundary conditions for general elastic collision models. The derivation of R13 equations is given in section 3, where we also demonstrate the super-Burnett order of the model. In section 4, we formulate the boundary conditions, analyze the boundary layers, and remove the unwanted layers by fixing the boundary conditions. One-dimensional examples are shown in section 5 to validate our models. We conclude our paper with a brief summary in section 6.
@@ 2.1. Choice of variables
The 13-moment equations were first proposed by [17], where the distribution function was expanded into an infinite series using Hermite polynomials. The coefficients of the polynomials were regarded as moments, and Grad chose to include the first 13 moments, namely the density ($\rho$), momentum ($\rho v_{i}$), pressure ($\rho\theta$), stress tensor ($\sigma_{ij}$) and heat flux ($q_{i}$), into his equations. In later developments of 13-moment systems, such a choice of the 13 moments becomes a standard and is followed by most works [34, 25, 44]. In particular, for the regularized 13-moment equations for Maxwell molecules, such a choice of moments matches exactly with the analysis using the order of magnitude method [32, Chapter 8]. It is seen that for linear equations, these moments, denoted by $\langle\boldsymbol{\phi}_{13},f\rangle$ with $\boldsymbol{\phi}_{13}$ being a 13-dimensional vector of polynomials of $\boldsymbol{\xi}$, fully describes the distribution function $f$ up to order $O(\mathrm{Kn})$, which means for any polynomial $r(\boldsymbol{\xi})$ satisfying $\langle\boldsymbol{\phi}_{13},r\rangle=0$, the moment $\langle r,f\rangle$ has at least order $O(\mathrm{Kn}^{2})$ when performing asymptotic analysis. This property leads to the super-Burnett order of R13 equations for Maxwell molecules, meaning that the super-Burnett equations can be derived from the R13 equations by Chapman-Enskog expansion. However, it is also known that this no longer holds for other types of molecules [31]. Inspired by this fact, in our study of 13-moment equations for general gas molecules, we will adopt a different set of variables $\langle\tilde{\boldsymbol{\phi}}_{13},f\rangle$, such that we can again obtain $\langle r,f\rangle\sim o(\mathrm{Kn})$ once $\langle\tilde{\boldsymbol{\phi}}_{13},r\rangle=0$. This will help achieve the super-Burnett order for the moment equations.
@@ 2.2. Moment equations and boundary conditions
The first three equations (7)(8)(9) are conservation laws of mass, energy and momentum. By comparing (8) and (9) with the conservation laws expressed by the stress tensor and the heat flux:
@@ 5.1. Steady-state examples
To compare with the example in sections 4.2 & 4.3, we consider a steady-state problem here. In fact, equations (77)–(81) can be regarded as a portion of the considering one-dimensional steady-state channel flows by the following replacement: $x\to x_{2}$, $v\to v_{2}$, $\bar{q}\to\bar{q}_{2}$ and $\bar{\sigma}\to\bar{\sigma}_{22}$. The boundary conditions (82)–(85) also share the same form of the considering boundary conditions by the above replacement. We adopt the coefficients $\theta^{W}_{l}=0$ and $\theta^{W}_{r}=0.2$, consistent with the example presented in section 4.2.
@@ 6. Conclusion
This work provides a new perspective of the derivation of time-dependent regularized 13-moment equations from linear kinetic equation under general elastic collision models. Compared with previous works such as [19] and [10], our equations possess not only the super-Burnett order, but also a symmetric structure that is useful for deriving Onsager boundary conditions. Another contribution of this work is the technique to remove undesired boundary layers, which needs to be applied since these layers are obviously unphysical.

### v3

@@ 1. Introduction
Instead of directly solving the Boltzmann equation, many researchers focus on the macroscopic moments of the distribution function and modify classical fluid equations, such as the Navier-Stokes equations, to describe the mechanics of moderately rarefied gases. Two classical techniques used to derive higher-order (in terms of Knudsen number) macroscopic fluid equations from the Boltzmann equation are the Grad’s moment method in [18] and the Chapman-Enskog expansion in [13]. Despite their successes in capturing near-equilibrium effects, the Grad’s 13-moment equations may experience a loss of hyperbolicity and unphysical subshocks. Moreover, the second-order Burnett and third-order super-Burnett equations, which are derived from the Chapman-Enskog expansion, suffer from instability. Recent researches address these limitations by employing various regularization techniques ([23, 26, 36, 7, 28, 19]) to enhance the accuracy and stability of the methods across a broader range of rarefied gas dynamics. Besides the regularization by the above two classical techniques, researchers have also explored alternative ways to address their limitations and derive higher-order equations such as the Onsager-consistent approaches ([1, 21]).
@@
In this work, we will investigate the regularized version of Grad’s 13-moment equations proposed in [36], which can be interpreted as a Chapman–Enskog like procedure applied to the Grad’s 13 moments. This regularization, abbreviated as R13, has been developed for both Maxwell ([34]) and non-Maxwell molecules ([38, 12]) and validated across diverse applications ([39, 30, 41]). The R13 equations are attractive since the equations include only second-order derivatives while attaining the super-Burnett order. Here the super-Burnett order means that the super-Burnett equations, which are the result of the Chapman-Enskog expansion up to the third order, can be derived from the R13 equations. Note that the original super-Burnett equations contain fourth-order derivatives, causing significant difficulties in the numerical discretization and formulation of boundary conditions. Due to the inclusion of corrections from higher-order moments through terms derived from the Chapman-Enskog expansion, the R13 equations can describe various rarefaction effects such as Knudsen boundary layers and heat fluxes from cold areas to hot areas ([29]). Also, the R13 equations are equipped with reliable boundary conditions ([43, 35]) and have been shown to be robust in numerical simulations ([29]). Recently in [42], Onsager boundary conditions are proposed to ensure the stability of boundary value problems of linear R13 equations for Maxwell molecules. However, it is not straightforward to generalize these Onsager boundary conditions to non-Maxwell molecules, even in the linear case. The main reason is the loss of a symmetric structure required by the Onsager boundary conditions (see [8]). To remedy this, in [11], the linear R13 equations for arbitrary elastic collision models were re-derived using a different method. The new derivation relies on the simpler form of the Chapman-Enskog expansion for steady-state equations compared with the time-dependent case, so that it cannot be applied directly to dynamic problems.
@@
The rest of this paper is organized as follows. Our linearization and nondimensionalization of the Boltzmann equation and moment equations are introduced in Section 2. Section 3 introduces the variables and then shows our main results including the explicit expressions of linear R13 equations with boundary conditions for general elastic collision models. The derivation of R13 equations is given in section 4, where we also demonstrate the super-Burnett order of the model. In section 5, we formulate the boundary conditions, analyze the boundary layers, and remove the unwanted layers by fixing the boundary conditions. One-dimensional examples are shown in section 6 to validate our models. We conclude our paper with a brief summary in section 7.
2. Linearization and nondimensionalization
All the derivations in this work will be established on dimensionless and linearized equations. In this section, we will state our approaches to both linearization and nondimensionalization. Note that each dimensionless variable in this section will be labeled by adding a hat “$\,\hat{\cdot}\,$” on top of the symbol denoting the variable, while such accents will be removed for convenience in other sections.
2.1. Linearization and nondimensionalization of the Boltzmann equation
In the original Boltzmann equation, the distribution function $f(\boldsymbol{x},\boldsymbol{\xi},t)$ satisfies the following integro-differential equation:
(7) $\frac{\partial f}{\partial t}+\xi_{k}\frac{\partial f}{\partial x_{k}}=Q[f,f]:=\int_{\mathbb{R}^{3}}\int_{\mathbb{S}^{2}}B(\boldsymbol{\xi}-\boldsymbol{\xi}_{1},\boldsymbol{\omega})[f(\boldsymbol{\xi}_{1}^{\prime})f(\boldsymbol{\xi}^{\prime})-f(\boldsymbol{\xi}_{1})f(\boldsymbol{\xi})]\,\mathrm{d}\boldsymbol{\omega}\,\mathrm{d}\boldsymbol{\xi}_{1},$
where $B(\cdot,\cdot)$ stands for the collision kernel, and
$\boldsymbol{\xi}^{\prime}=\frac{\boldsymbol{\xi}+\boldsymbol{\xi}_{1}}{2}+\frac{|\boldsymbol{\xi}-\boldsymbol{\xi}_{1}|}{2}\boldsymbol{\omega},\qquad\boldsymbol{\xi}_{1}^{\prime}=\frac{\boldsymbol{\xi}+\boldsymbol{\xi}_{1}}{2}-\frac{|\boldsymbol{\xi}-\boldsymbol{\xi}_{1}|}{2}\boldsymbol{\omega}.$
In this work, we are interested in the linear regime where the distribution function is close to a global Maxwellian with density $\rho_{0}$, velocity $\boldsymbol{v}_{0}$, temperature $\theta_{0}$ and mass of a single molecule $m$:
$\mathcal{M}_{0}(\boldsymbol{\xi})=\frac{\rho_{0}}{m(2\pi\theta_{0})^{3/2}}\exp\left(-\frac{|\boldsymbol{\xi}-\boldsymbol{v}_{0}|^{2}}{2\theta_{0}}\right).$
We can then introduce a small quantity $\epsilon$ and write the distribution function $f$ as
(8) $f(\boldsymbol{x},\boldsymbol{\xi},t)=\mathcal{M}_{0}(\boldsymbol{\xi})[1+\epsilon\hat{f}(\hat{\boldsymbol{x}},\hat{\boldsymbol{\xi}},\hat{t})],$
where $\hat{\boldsymbol{x}}$, $\hat{\boldsymbol{\xi}}$ and $\hat{t}$ are dimensionless variables defined by
(9) $\hat{\boldsymbol{x}}=(\boldsymbol{x}-\boldsymbol{v}_{0}t)/L,\qquad\hat{\boldsymbol{\xi}}=(\boldsymbol{\xi}-\boldsymbol{v}_{0})/\sqrt{\theta_{0}},\qquad\hat{t}=t/(L/\sqrt{\theta_{0}}).$
Plugging (8) into the Boltzmann equation yields
(10) $\frac{\partial\hat{f}}{\partial\hat{t}}+\hat{\xi}_{k}\frac{\partial\hat{f}}{\partial\hat{x}_{k}}=\frac{\rho_{0}L}{m\sqrt{\theta_{0}}}\frac{\mathcal{Q}[\hat{\mathcal{M}}_{0},\hat{f}\hat{\mathcal{M}}_{0}]+\mathcal{Q}[\hat{f}\hat{\mathcal{M}}_{0},\hat{\mathcal{M}}_{0}]+\epsilon\mathcal{Q}[\hat{f}\hat{\mathcal{M}}_{0},\hat{f}\hat{\mathcal{M}}_{0}]}{\hat{\mathcal{M}}_{0}},$
where ^$\hat{\mathcal{M}}_{0}$ is the dimensionless Maxwellian
^$\hat{\mathcal{M}}_{0}(\hat{\boldsymbol{\xi}})=\frac{1}{(2\pi)^{3/2}}\exp\left(-\frac{|\hat{\boldsymbol{\xi}}|^{2}}{2}\right).$
The linearized equation can then be obtained by discarding the $O(\epsilon)$ term in (10) and defining
^$\hat{\mathcal{L}}[\hat{f}]:=\frac{\mathcal{Q}[\hat{\mathcal{M}}_{0},\hat{f}\hat{\mathcal{M}}_{0}]+\mathcal{Q}[\hat{f}\hat{\mathcal{M}}_{0},\hat{\mathcal{M}}_{0}]}{B_{0}\hat{\mathcal{M}}_{0}},\qquad\mathrm{Kn}=\frac{mB_{0}\sqrt{\theta_{0}}}{\rho_{0}L}.$
Here $B_{0}$ is a constant with the same dimension as the collision kernel $B(\cdot,\cdot)$ to guarantee that $\hat{\mathcal{L}}$ is dimensionless, and here we choose $B_{0}$ such that
$\langle\hat{\xi}_{1}\hat{\xi}_{2},\hat{\mathcal{L}}(\hat{\xi}_{1}\hat{\xi}_{2})\rangle=-1,$
where the inner product is defined by
$\langle\hat{f},\hat{g}\rangle=\int_{\mathbb{R}^{3}}\hat{f}(\hat{\boldsymbol{\xi}})\hat{g}(\hat{\boldsymbol{\xi}})\hat{\mathcal{M}}_{0}(\hat{\boldsymbol{\xi}})\,\mathrm{d}\hat{\boldsymbol{\xi}}.$
Note that the linearized Boltzmann equation satisfies the linearized H-theorem. In the original Boltzmann equation (7), the entropy density and the entropy flux are defined by
$\eta=-k_{B}\int_{\mathbb{R}^{3}}f\log f\,\mathrm{d}\boldsymbol{\xi},\qquad\psi_{k}=-k_{B}\int_{\mathbb{R}^{3}}\xi_{k}f\log f\,\mathrm{d}\boldsymbol{\xi},$
which satisfy the inequality
$\frac{\partial\eta}{\partial t}+\frac{\partial\psi_{k}}{\partial x_{k}}\geqslant 0,$
indicating the second law of thermodynamics. For the linearized Boltzmann equation, the corresponding definitions are
$\hat{\eta}=-\int_{\mathbb{R}^{3}}|\hat{f}|^{2}\,\mathrm{d}\hat{\boldsymbol{\xi}},\qquad\hat{\psi}_{k}=-\int_{\mathbb{R}^{3}}\xi_{k}|\hat{f}|^{2}\,\mathrm{d}\hat{\boldsymbol{\xi}}.$
Since the linearized collision operator $\hat{\mathcal{L}}$ satisfies $\langle\hat{f},\hat{\mathcal{L}}[\hat{f}]\rangle\leqslant 0$ for any $\hat{f}$, the H-theorem turns out to be the $L^{2}$ stability of the linearized Boltzmann equation:
$\frac{\partial\hat{\eta}}{\partial\hat{t}}+\frac{\partial\hat{\psi}_{k}}{\partial\hat{x}_{k}}\geqslant 0,$
Later in our derivation of moment equations, this property will be preserved.
2.2. Linearization and nondimensionalization of the conservation laws
The laws of mass, momentum and energy conservation can be derived by taking moments of the Boltzmann equation, which yields
$\frac{\partial\rho}{\partial t}+\frac{\partial(\rho v_{j})}{\partial x_{j}}=0,$
$\frac{\partial(\rho v_{i})}{\partial t}+\frac{\partial}{\partial x_{j}}(\rho v_{i}v_{j}+\rho\theta\delta_{ij}+\sigma_{ij})=0,$
$\frac{\partial}{\partial t}\left(\frac{1}{2}\rho v_{i}v_{i}+\frac{3}{2}\rho\theta\right)+\frac{\partial}{\partial x_{j}}\left(\frac{5}{2}\rho\theta v_{j}+\frac{1}{2}\rho v_{i}v_{i}v_{j}+\sigma_{ij}v_{i}+q_{j}\right)=0.$
Here $\sigma_{ij}$ denotes the stress tensor and $q_{j}$ denotes the heat flux. To linearize these equations, we again assume that the fluid state is close to the equilibrium specified by density $\rho_{0}$, velocity $v_{0,i}$ and temperature $\theta_{0}$, and introduce dimensionless variables $\hat{\rho}$, $\hat{\theta}$, $\hat{v}_{i}$, $\hat{\sigma}_{ij}$ and $\hat{q}_{i}$ by
$\rho=\rho_{0}(1+\epsilon\hat{\rho}),\quad\theta=\theta_{0}(1+\epsilon\hat{\theta}),\quad v_{i}=v_{0,i}+\epsilon\sqrt{\theta_{0}}\hat{v}_{i},\quad\sigma_{ij}=\epsilon\rho_{0}\theta_{0}\hat{\sigma}_{ij},\quad q_{i}=\epsilon\rho_{0}\theta_{0}^{3/2}\hat{q}_{i}.$
By substituting the above expressions together with (9) into the conservation laws, and then dropping all the higher-order terms, we reach the linearized and dimensionless conservation laws:
(11) $\frac{\partial\hat{\rho}}{\partial\hat{t}}+\frac{\partial\hat{v}_{j}}{\partial\hat{x}_{j}}=0,$
$\frac{\partial\hat{v}_{i}}{\partial\hat{t}}+\frac{\partial\hat{\rho}}{\partial\hat{x}_{i}}+\frac{\partial\hat{\theta}}{\partial\hat{x}_{i}}+\frac{\partial\hat{\sigma}_{ij}}{\partial\hat{x}_{j}}=0,$
$\frac{\partial\hat{\theta}}{\partial\hat{t}}+\frac{2}{3}\frac{\partial\hat{v}_{j}}{\partial\hat{x}_{j}}+\frac{2}{3}\frac{\partial\hat{q}_{j}}{\partial\hat{x}_{j}}=0.$
Another approach to derive the same equations is to directly take moments of the linearized Boltzmann equation. The relationship between the dimensionless moments ($\hat{\rho}$, $\hat{\theta}$, etc.) and the dimensionless distribution function $\hat{f}$ is
$\hat{\rho}=\langle 1,\hat{f}\rangle,\quad\hat{v}_{i}=\langle\hat{\xi}_{i},\hat{f}\rangle,\quad\hat{\theta}=\left\langle\frac{1}{3}|\hat{\boldsymbol{\xi}}|^{2}-1,\hat{f}\right\rangle,$
$\hat{\sigma}_{ij}=\left\langle\hat{\xi}_{i}\hat{\xi}_{j}-\frac{1}{3}|\hat{\boldsymbol{\xi}}|^{2}\delta_{ij},\hat{f}\right\rangle,\qquad\hat{q}_{i}=\left\langle\frac{1}{3}|\hat{\boldsymbol{\xi}}|^{2}\hat{\xi}_{i},\hat{f}\right\rangle.$
Thus, using the fact that $\langle 1,\hat{\mathcal{L}}[\hat{f}]\rangle=\langle\hat{\xi}_{i},\hat{\mathcal{L}}[\hat{f}]\rangle=\langle|\hat{\boldsymbol{\xi}}|^{2},\hat{\mathcal{L}}[\hat{f}]\rangle=0$, one can take the inner product of the linearized equation (1) and the polynomials $1$, $\hat{\xi}_{i}$ and $|\hat{\boldsymbol{\xi}}|^{2}$ to obtain the moment equations (11).
The two methods to formulate the conservation laws are essentially equivalent, since taking moments and linearizing equations are two commutable operations. In our derivation of moment equations to be presented below, we will follow the second approach which takes moments of the linearized equations (1). Due to the presence of $\hat{\sigma}_{ij}$ and $\hat{q}_{i}$ in (11), our purpose is to find a suitable closure that provides decent approximation to the linearized Boltzmann equation when the Knudsen number $\mathrm{Kn}$ is small. Hereafter, the accent “ $\hat{\cdot}$ ” will no longer be added to dimensionless variables.
@@ 3.1. Choice of variables
The 13-moment equations were first proposed by [18], where the distribution function was expanded into an infinite series using Hermite polynomials. The coefficients of the polynomials were regarded as moments, and Grad chose to include the first 13 moments, namely the density ($\rho$), momentum ($\rho v_{i}$), pressure ($\rho\theta$), stress tensor ($\sigma_{ij}$) and heat flux ($q_{i}$), into his equations. In later developments of 13-moment systems, such a choice of the 13 moments becomes a standard and is followed by most works [36, 27, 46]. In particular, for the regularized 13-moment equations for Maxwell molecules, such a choice of moments matches exactly with the analysis using the order of magnitude method [34, Chapter 8]. It is seen that for linear equations, these moments, denoted by $\langle\boldsymbol{\phi}_{13},f\rangle$ with $\boldsymbol{\phi}_{13}$ being a 13-dimensional vector of polynomials of $\boldsymbol{\xi}$, fully describes the distribution function $f$ up to order $O(\mathrm{Kn})$, which means for any polynomial $r(\boldsymbol{\xi})$ satisfying $\langle\boldsymbol{\phi}_{13},r\rangle=0$, the moment $\langle r,f\rangle$ has at least order $O(\mathrm{Kn}^{2})$ when performing asymptotic analysis. This property leads to the super-Burnett order of R13 equations for Maxwell molecules, meaning that the super-Burnett equations can be derived from the R13 equations by Chapman-Enskog expansion. However, it is also known that this no longer holds for other types of molecules [33]. Inspired by this fact, in our study of 13-moment equations for general gas molecules, we will adopt a different set of variables $\langle\tilde{\boldsymbol{\phi}}_{13},f\rangle$, such that we can again obtain $\langle r,f\rangle\sim o(\mathrm{Kn})$ once $\langle\tilde{\boldsymbol{\phi}}_{13},r\rangle=0$. This will help achieve the super-Burnett order for the moment equations.
@@ 3.2. Moment equations and boundary conditions
The first three equations (12)(13)(14) are conservation laws of mass, energy and momentum. By comparing (13) and (14) with the conservation laws expressed by the stress tensor and the heat flux (see (11)):
@@ 6.1. Steady-state examples
To compare with the example in sections 5.2 & 5.3, we consider a steady-state problem here. In fact, equations (82)–(86) can be regarded as a portion of the considering one-dimensional steady-state channel flows by the following replacement: $x\to x_{2}$, $v\to v_{2}$, $\bar{q}\to\bar{q}_{2}$ and $\bar{\sigma}\to\bar{\sigma}_{22}$. The boundary conditions (87)–(90) also share the same form of the considering boundary conditions by the above replacement. We adopt the coefficients $\theta^{W}_{l}=0$ and $\theta^{W}_{r}=0.2$, consistent with the example presented in section 5.2.
@@ 7. Conclusion
This work provides a new perspective of the derivation of time-dependent regularized 13-moment equations from linear kinetic equation under general elastic collision models. Compared with previous works such as [20] and [11], our equations possess not only the super-Burnett order, but also a symmetric structure that is useful for deriving Onsager boundary conditions. In the linear regime, our equations can well approximate the Boltzmann equation with moderate Knudsen numbers and preserve the conservation laws and the dissipation of entropy. Another contribution of this work is the technique to remove undesired boundary layers, which needs to be applied since these layers are obviously unphysical.
@@
Declaration of interests.
The author reports no conflict of interest.